# Enhancing multimodal analogical reasoning with Logic Augmented Generation

### v1

@@ III-A Background
Situations can be reported from different perspectives: the Cognitive Perspectivisation Ontology  [46] represents the Perspectivisation frame, a type of situations where a fact is reported within a certain narrative, or Lens, which creates a viewpoint, cutting the reality which a Conceptualiser holds an attitude towards. Perspectivisation itself is therefore not merely rhetorical, as it blends entities playing two roles: the cut and the lens. Finally, the Amnestic Forgery ontology  [47] is a model for metaphor semantics based on MetaNet and compatible with Conceptual Metaphor Theory [49]. Amnestic Forgery reuses and extends the Framester schema, allowing to deal with both semiotic and referential aspects of frames.
III-B Ontology
@@
The proposed ontology generalises the notion of metaphors as frame mappings, and uses mDnS for ground axiomatisations. The ontology defines key classes and properties that operationalize the blending of concepts:
@@
    Blended, the description that inherits components from the Blendable frames, firstly through the Blending sharable space, and eventually inheriting other components from the Blendable frames, giving rise to a novel set of blended situations;
@@
Figure 3: The core of the Blending ontology [^9] The full diagram can be seen on Github https://github.com/dersuchendee/knowledge-augmented-generation.
@@
IV Methodology
The proposed method, based on LAG [22], aims to enhance the detection and understanding of metaphor by establishing a standardized and explainable model for the semantic representation of implicit natural language. Therefore, our approach has two main components: i) the representation of inputs, regardless of their modality, into natural language, ii) their conversion into a base knowledge graph with Text2AMR2FRED and iii) the SKG enhancement driven by Conceptual Blending Theories applied to metaphor.
IV-A Automated natural language and image KG representation with Text2AMR2FRED
Figure 4: Text2AMR2FRED pipeline for the sentence “Crime has infected communities everywhere”.
AMR identifies the cause of the infection in the sentence, which is, however, not a germ, but crime and the thing that gets infected (community) everywhere. In the FRED-like KG, this graph is enriched with knowledge from Framester such as frames, WordNet synsets and PropBank roles, along with alignment to the DOLCE foundational ontology  [52]. However, there is no representation of metaphorical knowledge yet at this stage. While a conceptual metaphor KG sourced from Metanet is contained in Framester, it is still necessary to be able to detect figurative language in any text. An image can also be used by being automatically described in natural language. In this case, language is an amodal way to bridge modalities in the resulting supramodal SKG. The SKG can then be extended with implicit contextual knowledge with the support of an LLM.
IV-B Logic Augmented Generation with heuristics
Figure 5: Logic augmented graph for the sentence “Crime has infected communities everywhere”. In black is the base graph, in red the augmented classes and properties.
As shown in Fig. 5, the result of the framework that generates XKG triples is the merging of the original KG with new classes and properties according to a specific ontology in the prompt. In this case, a minimum requirement of selected classes of the Blending Ontology is specified to account for the combination of concepts. Considering this framework, it is possible to use Conceptual Blending specifically for metaphor, where the two domains, the source and the target, correspond to the Blendable frames. These frames have an overarching Blending frame in common that both explains and motivates the metaphor. Such fine-grained element mappings are required to understand the consistency of the blend, something that has been mentioned in research but so far has yet to be implemented [40]. The Blending frame motivating the metaphorical blend also makes a system’s analogical reasoning capabilities more explainable. For example, a system can correctly identify the elements involved in a metaphor but misidentify the property that links them, as discussed in Section VII, showing the inability to understand the metaphor. With this framework in mind, it is possible to create a prompt for an LLM that, given a text and the resulting KG representation from Text2AMR2FRED, it extends it to create an augmented RDF graph in Turtle format.
@@ V Application and Evaluation
In this section, we evaluate the framework’s ability to extract implicit meaning by applying it to metaphor detection and understanding. First, we test the LAG framework on three metaphor detection datasets. The best configuration is then applied to three metaphor understanding datasets. Finally, we perform ablation tests on metaphor detection and visual metaphor understanding.
@@ V-A Tasks
This section shows the application of the framework in three metaphor-related tasks: i) metaphor detection, where a SKG derived from a sentence in natural language is enhanced to determine if a sentence is metaphorical, sometimes with a target word as a reference; ii) conceptual metaphor understanding, which additionally aims to identify the correct source and target domains that make up the metaphor, and iii) visual metaphor understanding, where the aim is to identify the property linking the two domains in the visual metaphor.
@@ V-A1 Metaphor detection and understanding
As shown in the Supplementary Material, the LLM is given a sentence, a SKG representing the sentence derived from Text2AMR2FRED, and a series of instructions. The prompt first requires identifying the possible source and target of the sentence; then, it is asked to define these two as conceptual frames containing roles and following a blending function that orchestrates the creation of a blended space. Following the Blending Ontology, a Lens and an Attitude further refine the definition of the metaphor. An example of the graph for a specific sentence is also provided.
Given the Text2AMR2FRED graph, and the presumed source and target of the sentence given a target word when available, along with blending-derived instructions, the LLM is prompted to analyze the sentence and to use at least these specific ontology elements according to the given definitions, to determine the possible metaphoricity of a sentence:
  1. 1.
    bl:Blending,
  2. 2.
    bl:Blendable,
  3. 3.
    bl:Blended,
  4. 4.
    cp:Attitude,
  5. 5.
    cp:Lens,
  6. 6.
    the datatype property metanet:isMetaphorical, which classification is based on.
Furthermore, in the prompt it is asked to determine blendable roles and their mapping to create a blended metaphorical space according to the Blending Ontology.
@@ V-A2 Visual metaphor understanding
Following the original study performed on human participants outlined in Petridis et al.  [32], the methodology for visual metaphor understanding involves the presentation of three examples and correct annotations of visual metaphor elements within an in-context learning technique, along with a new image to analyze. In this case, the image to analyze is first described in natural language, and the text is passed to Text2AMR2FRED to generate a base graph. The base graph, image, accompanying sentence, and blending instructions are all included in the prompt to generate the augmented KG, as shown in Fig. 6. In this work, we experiment with multiple combinations of these inputs to find the best prompting technique.
@@ VI Datasets
For the purpose of this study, and for comparison across the outlined tasks, for textual metaphor detection we conducted experiments on two publicly available metaphor datasets: (1) MOH-X [53], which comprises 647 sentences annotated as either metaphorical or literal; and (2) TroFi [54], another metaphor detection dataset collected from the 1987-1989 Wall Street Journal Corpus. We sampled 300 balanced instances each for testing. For conceptual metaphors, we employ (3) the dataset refined by Wachowiak and Gromann [40] of 447 sentences including conceptual metaphors annotated with source and target domain (for ease of reference, we will call it WG). We also introduced a fourth dataset to verify the capability to identify and explain domain-specific conceptual metaphors: (4) the Balanced Conceptual Metaphor Testing Dataset (BCMTD). This dataset consists of 147 sentences, equally distributed among general conceptual metaphors (derived from annotated Metanet-derived data from Framester [35] and that weren’t present in WG), scientific conceptual metaphors (extracted from scientific literature from the Appendix in Van Rijn and Van Tongeren [55] and examples from Semino et al. [56], allowing a domain-specific dataset on medical metaphors), and literal sentences from the VUA corpus [57] (previously refined by Wachowiak and Gromann [40] as many of the sentences were wrongly labeled or deemed unfit for the task). The visual metaphors dataset is derived from Petridis et al. [32]. It contains 48 images, 27 ads and 21 non-ads, and three examples, that have been given to human participants to test their ability to identify visual metaphors according to a conceptual blending framework. Table I shows statistics for all the datasets.
@@
BCMTD 170 (conceptual) 451 (scientific) 50 (VUA) 66.6 147
@@
VI-A Evaluation metrics
The evaluation metrics for metaphor detection are both automated and manual depending on the type of task.
@@ VI-A1 Automated evaluation metrics
We use Accuracy and F1 score to evaluate the model’s performance both in generic and conceptual metaphor detection. For conceptual metaphor understanding, evaluated via the correct identification of both source and target mappings according to a gold standard, the BLEURT score [58] has been shown to capture semantic similarities between texts. It is considered in order to evaluate the distance between the candidate and reference source and target in the metaphor. This is an evaluation method similar to the one in Saakyan et al. [59], where instead of only reporting label accuracy, it is reported according to defined thresholds. In this case, the threshold is where the correctly predicted labels of source and target domains achieve a BLEURT positive score.
@@ VI-A2 Manual evaluation
To correctly determine metaphorical source and target mappings for domain identification, and to correctly compare gold standard and predicted blending properties for visual metaphor interpretation, it is also necessary to conduct a manual evaluation. The best performing solution for conceptual and visual metaphor understanding is therefore evaluated for source and target mappings by three annotators. The annotators were graduated in linguistics, had proven expertise on metaphors, and had English Native or near-native language proficiency. They exhibited a Fair inter-annotator agreement (Fleiss’ Kappa) in all cases, with an average on all tasks of 0.35, which is understandable for a task so subjective as metaphor understanding. More information on the agreement scores is in the Supplementary Materials. To calculate the Accuracy, we compute an average of each Accuracy score. In the visual metaphor interpretation task, the manual evaluation compares the gold standard property with the predicted one connecting source and target. The evaluation directly follows the one in Petridis et al. [32], which considered as correct responses directly related to the objects in the image. We also repeated the study on a sample of 24 images on 20 participants of the University of Bologna, manually evaluating their responses as well.
@@ VI-B Baseline methods
In Tian et al. [37], several baselines for metaphor detection are discussed, with the best-performing method being a Theory Guided Scaffolding Instruction Framework based on Conceptual Metaphor Theory (TSI CMT). MetaPRO is a baseline among non-LLM methods for detecting source and target domains, serving as the state-of-the-art approach for identifying mappings without specifying a target word [36]. For conceptual metaphor understanding, no current baselines simultaneously detect source and target but only one of the two. We thus employ LLMs in different configurations, including zero-shot and few-shot, shown to be effective in previous research [40]. No established baselines for visual metaphor understanding specify source, target, and blending properties. Thus, we explore different configurations of the LAG framework using the same three few-shot examples used to test humans in [32], and evaluate against the few-shot baseline without LAG.
@@ VI-C Parameters setup
In our experiments, Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet. serves as the primary LLM. We compare the best-performing results with Llama 3.1 70B Instruct Turbo [^11] https://ai.meta.com/blog/meta-llama-3-1/ in all cases except for visual metaphor interpretation, where no current open models support few-shot image examples. Only the APIs are used for testing. We conducted tests using default parameters and set the temperature to 0 for greedy decoding.
@@ VI-D2 Conceptual metaphor understanding
Conceptual metaphor understanding is a more challenging task than simple metaphor detection, as it involves the correct detection of both source and target domains for a metaphor, which shows whether the system has understood it or not. We employ majority voting between the three annotators to select the gold standard label.
WG dataset. Manual evaluation performed by the three annotators for the dataset reveals that the percentage of metaphors where source and target were recorded to be correctly identified at the same time is 25.6%. BLEURT scores showed a moderate correlation between the human annotations and the predicted domains. For the target domain, a point-biserial correlation of 0.702 (p <<< 0.05) and a Spearman correlation of 0.698 indicate that BLEURT is reasonably sensitive to the human judgments. For the source domain, a point-biserial correlation of 0.707 (p <<< 0.05) and a Spearman correlation of 0.680 were observed.
BCTMD. Manual evaluation performed by the three annotators for the best performing methodology in metaphor understanding, in fair agreement, reveals that out of 96 conceptual metaphors, the percentage of correctly identified source and target at the same time is 34.3%. Furthermore, the distribution of correctly identified metaphors is 51.6% for generic conceptual metaphors, while for scientific conceptual metaphors the score lowers to 8%. This score, with respect to the generic metaphors also contained in WG, might be related to the fact that there is no domain-specific metaphor dataset on which LLMs have been trained, but also that scientific metaphors have different levels of conventionality, tending to result more literal as they come to be widespread in language. BLEURT scores showed a 36.9% positive correlation between reference and predicted source and target domains. A point-biserial correlation of 0.3525 (p <<< 0.05) and Spearman correlation of 0.2917 (p <<< 0.05) both indicate moderate positive correlations for BLEURT.
@@ VI-D3 Visual metaphor understanding
For visual metaphor understanding, the identification of the blending property corresponds to the correct interpretation of the source and target employed in the metaphor (as shown by Petridis et al.[32]). Out of the three configurations, shown in Table V, manual annotation from the three evaluators revealed that the best performing result is LAG without the insertion of a sentence describing the image alongside the image itself in the prompt. Concerning human evaluation described in Petridis et al. [32], which correctly interpreted the meaning of the image 41.32% of the time, in the best performing method the LLM correctly interpreted the meaning an average of 67.06% of the time. We then sampled the dataset, randomly taking 24 visual metaphors out of the 48 in the dataset, and repeated the original study by [32] with 20 human participants recruited from the University of Bologna, from student to researcher level, with native or near-native English fluency. This setup slightly differs from the original study, where participants were employed without specifying their study level. Participants’ submissions were anynomous and informed consent was provided. This time as well, on average, human participants have correctly interpreted the meaning 59.25% of the time, with respect to the average LLM score of 73.5% for the same images.

### v2

@@ IV-B1 The Blending Ontology
The Blending Ontology [^5] Available at https://github.com/dersuchendee/BlendingOntology is based upon mDnS (mereological Descriptions and Situations) [^6] http://www.ontologydesignpatterns.org/ont/mdns/mdns.owl, the Cognitive Perspectivisation ontology as defined in Gangemi and Presutti [59] [^7] http://www.ontologydesignpatterns.org/ont/persp/perspectivisation.owl, and the MetaNet schema from the Amnestic Forgery ontology by Gangemi et al. [60]. By introducing the concepts of descriptions and situations, the Description and Situations ontology design pattern [61] provides a formalization of frame semantics where a situation is a set of facts as described by an observer that is interpreted by a description, a schema that defines the entities observed in a situation. Situations can be reported from different perspectives: the Cognitive Perspectivisation Ontology proposed by Gangemi and Presutti [59] represents the Perspectivisation frame, a type of situations where a fact is reported within a certain narrative, or Lens, which creates a viewpoint, cutting the reality which a Conceptualiser holds an attitude towards. Perspectivisation itself is therefore not merely rhetorical, as it blends entities playing two roles: the cut and the lens. Finally, the Amnestic Forgery ontology  [60] is a model for metaphor semantics based on MetaNet and compatible with Conceptual Metaphor Theory [62]. Amnestic Forgery reuses and extends the Framester schema, allowing to deal with both semiotic and referential aspects of frames.
@@
The Blending Ontology defines key classes and properties that operationalize the blending of concepts:
@@
    Blended, the description that inherits components from the Blendable frames, firstly through the Blending sharable space, and eventually inheriting other components from the Blendable frames, giving rise to a novel set of blended situations. As a consequence, the Blended is considered as a novel entity, separately from the Blendable;
@@
Figure 3: The core of the Blending ontology [^9] The full diagram can be seen on Github https://github.com/dersuchendee/metaphorical-logic-augmented-generation.
@@
Such fine-grained element mappings are required to understand the consistency of the blend, something that has been mentioned in research but so far has yet to be implemented [35]. The Blending frame motivating the metaphorical blend also makes a system’s analogical reasoning capabilities more explainable. For example, a system can correctly identify the elements involved in a metaphor but misidentify the property that links them, as discussed in Section VII, showing the inability to understand the metaphor. With this framework in mind, it is possible to create a prompt for an LLM that, given a text and the resulting KG representation from Text2AMR2FRED, it extends it to create an augmented RDF graph in Turtle format.
@@ V Application and Evaluation
In this section, we evaluate the LAG-based framework’s ability to extract implicit meaning by applying it to metaphor detection and understanding tasks. We first use three metaphor detection datasets to determine the best-performing configuration among several baselines. This optimal configuration is then applied to three metaphor understanding datasets.
@@ V-A Tasks
This section describes the application of the framework in three metaphor-related tasks: i) metaphor detection, where a SKG derived from a sentence in natural language is enhanced to determine if a sentence is metaphorical, sometimes with a target word as a reference; ii) conceptual metaphor understanding, which additionally aims to identify the correct source and target domains that make up the metaphor, and iii) visual metaphor understanding, where the aim is to identify the property linking the two domains in the visual metaphor.
@@ V-A1 Metaphor detection and understanding
As shown in the Supplementary Material, in this task, the LLM is given a sentence, a SKG representing the sentence derived from Text2AMR2FRED, and a series of instructions. The prompt first requires identifying the possible source and target of the sentence; then, it is asked to define these two as conceptual frames containing roles and following a blending function that orchestrates the creation of a blended space. Following the Blending Ontology’s alignment to the Cognitive Perspectivization one (See Section IV), a Lens and an Attitude further refine the definition of the metaphor. An example of the output graph for a specific sentence is also provided.
Given the Text2AMR2FRED graph, and the presumed source and target of the sentence given a target word when available, along with blending-derived instructions, the LLM is prompted to analyze the sentence and to use at least these specific ontology elements according to the given definitions, to determine the possible metaphoricity of a sentence: bl:Blending, bl:Blendable, bl:Blended, cp:Attitude, cp:Lens, and the datatype property metanet:isMetaphorical which the classification is based on. In the prompt, we also ask to determine blendable roles and their mapping to create a blended metaphorical space according to the Blending Ontology.
@@ V-A2 Visual metaphor understanding
Following the original study performed on human participants outlined in Petridis et al. [52], the methodology for visual metaphor understanding involves the presentation of three examples and correct annotations of visual metaphor elements within an in-context learning technique, along with a new image to analyze. In this case, each image to analyze is first described in natural language through the LLM, and the text is passed to Text2AMR2FRED to generate a SKG. The SKG, image, accompanying sentence, and Blending Ontology heuristics are all included in the prompt with the three examples to generate the XKG, as shown in Fig. 4. In this work, we experiment with multiple combinations of these inputs to find the best prompting technique.
@@ VI Datasets
For the purpose of this study, and for comparison across the outlined tasks, for textual metaphor detection we conducted experiments on two publicly available metaphor datasets: (1) MOH-X [63], which comprises 647 sentences annotated as either metaphorical or literal; and (2) TroFi [64], another metaphor detection dataset collected from the 1987-1989 Wall Street Journal Corpus. We sampled 300 balanced instances each for testing. For conceptual metaphors, we employ (3) the dataset refined by Wachowiak and Gromann [35] of 447 sentences including conceptual metaphors annotated with source and target domain (for ease of reference, we will name it WG). We also introduced a fourth dataset to verify the capability to identify and explain domain-specific conceptual metaphors: (4) the Balanced Conceptual Metaphor Testing Dataset (BCMTD). This dataset consists of 147 sentences, equally distributed among general conceptual metaphors (derived from annotated Metanet-derived data from Framester [56] and that weren’t present in WG), scientific conceptual metaphors (extracted from scientific literature from the Appendix in Van Rijn and Van Tongeren [65] and examples from Semino et al. [66], allowing a domain-specific dataset on medical metaphors), and literal sentences from the VUA corpus [67] (previously refined by Wachowiak and Gromann [35] as many of the sentences were wrongly labeled or deemed unfit for the task). The visual metaphors dataset is derived from Petridis et al. [52]. It contains 48 images, 27 ads and 21 non-ads, and three examples, that have been given to human participants to test their ability to identify visual metaphors according to a conceptual blending framework. Table I shows statistics for all the datasets.
@@
BCMTD 170 (conceptual) 48 (scientific) 50 (VUA) 66.6 147
@@
VI-A Evaluation methodology
The evaluation methodology for metaphor detection is both automated and manual depending on the type of task.
@@ VI-A1 Automated evaluation metrics
We use Accuracy and F1 score to evaluate the model’s performance in metaphor detection. For conceptual metaphor understanding, evaluated via the correct identification of both source and target mappings according to a gold standard, the BLEURT score [68] is used in order to evaluate the semantic distance between the candidate and reference source and target in the metaphor and considered to be successful if both achieve a positive score. This is an evaluation method similar to the one in Saakyan et al. [69], where instead of only reporting label accuracy, it is reported according to defined thresholds. We also compute the correlation between the BLEURT score and human evaluation.
@@ VI-A2 Manual evaluation
To correctly determine metaphorical source and target mappings for domain identification, and to correctly compare gold standard and predicted blending properties for visual metaphor interpretation, it is also necessary to conduct a manual evaluation. The best performing solution for conceptual and visual metaphor understanding is therefore evaluated for source and target mappings by three annotators. The annotators were graduated in linguistics, had proven expertise on metaphors, and had English Native or near-native language proficiency. They exhibited a Fair inter-annotator agreement (Fleiss’ Kappa) in all cases, with an average on all tasks of 0.35, which is understandable for a task so subjective as metaphor understanding. To select the gold standard label, we employ majority voting between the annotators. More information on the agreement scores is in the Supplementary Materials. To calculate the Accuracy, we compute an average of each Accuracy score. In the visual metaphor understanding task, the manual evaluation compares the gold standard property with the predicted one connecting source and target. The evaluation directly follows the one in Petridis et al. [52], which considered as correct responses directly related to the property in the image. We also repeated the study on a sample of 24 images on 20 participants of the University of Bologna, manually evaluating their responses as well. Participants ranged from student to researcher level, with native or near-native English fluency. This setup slightly differs from the original study, where participants were employed without specifying their study level. Participants’ submissions were anonymous and informed consent was provided. Furthermore, we provide a sample of 16 generated graphs, 8 derived from correctly identified metaphorical sentences, 8 derived from visual metaphors, to two knowledge engineers experts in conceptual blending and conceptual metaphor theories for an evaluation of the quality of the graphs. We evaluate the representation’s adequacy, correctness, and completeness on a Likert scale from 1 to 7 concerning detected property, source, target and element mappings. Finally, one knowledge engineer manually evaluated errors (derived from negative majority voting) for conceptual metaphor understanding of the best performing LAG configuration. The errors follow the categorization proposed in Wachowiak and Gromann [35] with the addition of a new category found, the incorrect switching of source and target (for a comprehensive account of these categories, see the Supplementary Material).
@@ VI-B Baseline methods
In Tian et al. [32] the best-performing method for metaphor detection is a Theory Guided Scaffolding Instruction Framework based on Conceptual Metaphor Theory (TSI CMT). MetaPRO is a baseline among non-LLM methods for detecting source and target domains, serving as the state-of-the-art approach for identifying mappings without specifying a target word [31]. For conceptual metaphor understanding, no current baselines simultaneously detect source and target but only one of the two. We thus employ LLMs in different configurations, including zero-shot and few-shot, shown to be effective in previous research [35]. No established baselines for visual metaphor understanding specify source, target, and blending properties. Thus, we explore different configurations of the LAG framework using the same three few-shot examples used to test humans in [52], and evaluate against the few-shot baseline without LAG.
@@ VI-C Parameters setup
In our experiments, Claude 3.5 Sonnet [^10] https://www.anthropic.com/news/claude-3-5-sonnet serves as the primary LLM. We compare the best-performing results with Llama 3.1 70B-Instruct [^11] https://ai.meta.com/blog/meta-llama-3-1/ in all cases except for visual metaphor interpretation, where no current open models support few-shot image examples. Only the APIs are used for testing. We conducted tests using default parameters and set the temperature to 0 for greedy decoding.
@@ VI-D2 Conceptual metaphor understanding
Conceptual metaphor understanding is a more challenging task than metaphor detection, as it involves the correct identification of both source and target domains for a metaphor, which shows whether the system has understood it or not.
WG dataset. Manual evaluation reveals that the percentage of metaphors where source and target were correctly identified at the same time is 25.6%. BLEURT scores showed a moderate correlation between the human annotations and the predicted domains. For the target domain, a point-biserial correlation of 0.702 (p $<$ 0.001) and a Spearman correlation of 0.698 indicate that BLEURT is reasonably sensitive to the human judgments. For the source domain, a point-biserial correlation of 0.707 (p $<$ 0.001) and a Spearman correlation of 0.680 were observed.
BCTMD. Manual evaluation on the best performing method in metaphor understanding reveals that out of 96 sentences containing conceptual metaphors, the percentage of correctly identified source and target at the same time is 34.3%. Furthermore, the distribution of correctly identified metaphors is 51.6% for generic conceptual metaphors, while for scientific conceptual metaphors the score lowers to 8%. BLEURT scores showed a 36.9% positive correlation between reference and predicted source and target domains. A point-biserial correlation of 0.3525 (p $<$ 0.001) and Spearman correlation of 0.2917 (p $<$ 0.001) both indicate moderate positive correlations for BLEURT.
@@ VI-D3 Visual metaphor understanding
For visual metaphor understanding, the identification of the blending property corresponds to the correct interpretation of the source and target employed in the metaphor (as shown by Petridis et al.[52]). Out of the three configurations, shown in Table V, manual annotation from the three evaluators revealed that the best performing result is LAG without the insertion of a sentence describing the image alongside the image itself in the prompt. Concerning human evaluation described in Petridis et al. [52], which correctly interpreted the meaning of the image 41.32% of the time, in the best performing method the LLM correctly interpreted the meaning an average of 67.06% of the time. We then sampled the dataset, randomly taking 24 visual metaphors out of the 48 in the dataset, and repeated the original study  [52] with 20 human participants. They have correctly interpreted the metaphorical property 59.25% of the time, with respect to the average LLM score of 73.5% for the same images.